OpenStudy (anonymous):

A triangle has one side that measures x, and the other two sides each measure 4 inches less than x. The perimeter is 19 inches. What is the measure of x?

OpenStudy (anonymous):

We know that \[P=\Sigma s = s _{1}+s_{2}+s_{3}=19\] We also know that \[s_{1}=x\] and \[s_{2}=s_{3}=x-4\] by substitution: \[19=x+(x-4)+(x-4)=3x-8\] so: \[27=3x\]\[x=9\] To check: \[9+(9-4)+(9-4)=19\]\[19=19\] CHECK

OpenStudy (anonymous):

so 9 right?